OpenStudy (anonymous):

when written with a rational denominator, the expression √2/(√2+√3-√5) is equivalent to?

OpenStudy (anonymous):

okay the first step is to multiply by its conjugate so \[\sqrt{2}\div \sqrt{2}+\sqrt{3} -\sqrt{5} \] times \[(\sqrt{2} +\sqrt{3} +\sqrt{5})\div \sqrt{2}+\sqrt{2} +\sqrt{5}\]

OpenStudy (anonymous):

and then simplify the radicals after you mutiply